Two balls are drawn in succession without replacement from an urn containing 4 red balls and 3 black balls. Find the probability distribution of the number of black balls. 


1
Expert's answer
2021-06-23T17:51:18-0400

Number of Red balls "=4"  

Number of Black balls "=3"

Total number of balls "=4+3=7"

Given: two balls are drawn in succession without replacement. 

Let "X=" the number of black balls: "X=0,1,2."


"P(X=0)=\dfrac{\dbinom{4}{2}\dbinom{3}{0}}{\dbinom{7}{2}}=\dfrac{6}{21}=\dfrac{2}{7}"


"P(X=1)=\dfrac{\dbinom{4}{1}\dbinom{3}{1}}{\dbinom{7}{2}}=\dfrac{4(3)}{21}=\dfrac{4}{7}"

"P(X=2)=\dfrac{\dbinom{4}{0}\dbinom{3}{2}}{\dbinom{7}{2}}=\dfrac{3}{21}=\dfrac{1}{7}"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 0 & 1 & 2 \\ \hline
 p(x) & 2/7 & 4/7 & 1/7 \\
\end{array}"

"E[X]=0(\dfrac{2}{7})+1(\dfrac{4}{7})+2(\dfrac{1}{7})=\dfrac{6}{7}"